OpenStudy (anonymous):

whats another way of writing \[\cos ^{2}\]?

OpenStudy (anonymous):

you can also write it \[(\cos(x))^2\]

OpenStudy (anonymous):

so i can apply chain rule to it?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

hey can I ask you another question?

OpenStudy (anonymous):

fire away

OpenStudy (anonymous):

how do I turn \[\sqrt{y}= x^{2}(x+7)\] into y=mx+b form?

OpenStudy (anonymous):

well first I would distribute the \[x^2\] into the (x+7) then solve for y by squaring both sides.

OpenStudy (anonymous):

my bad its : \[y^{2}= x^{2} (x+7)\]

OpenStudy (anonymous):

well then distribute the \[x2^\] then take the square root of both sides to solve for y. Remember when taking the square root you might need to account for the result being + or -

OpenStudy (anonymous):

err \[x^2\]